# Deep Reasoning Translation via Reinforcement Learning

### v1

@@ 3.1 Experimental Setups
Dataset. Following Wang et al. (2024), we use MetaphorTrans [^2] https://github.com/krystalan/DRT in our experiments, which is a English-to-Chinese literature translation dataset involving 19K training samples and 2K testing samples. Each sample involves a source English sentence, the corresponding thought and Chinese translation. The source sentences are selected from English literature books, and generally contain metaphors or similes, and thus tend to require deep thinking during translation. The thought and translations are synthesized by Qwen2.5-72B. In our experiments, we only use the source sentences of its training samples in our RL training framework.
In addition to the MetaphorTrans test set, we purchase electronic copies of two complete literature books, and use them to evaluate deep reasoning MT models: (1) The Essential O. Henry Collection (by O. Henry) and (2) Orbital (by Samantha Harvey). [^3] Due to copyright protection, we will only release the test data of The Essential O. Henry Collection and not release the content of Orbital (published in 2023).
Metrics. As mentioned by Wang et al. (2024), the references in MetaphorTrans test set are synthesized via LLMs, and are not ensured by human translators. Besides, the golden references of The Essential O. Henry Collection and Orbital are missing. Therefore, we adopt reference-free metrics in our experiments. Specifically, we use “CometKiwi” [^4] https://github.com/Unbabel/COMET Rei et al. (2022) to evaluate the model translations, which uses a language model to judge whether a model translation conveys the semantics of the source sentence. Moreover, following Wang et al. (2024), we use evaluators implemented using GPT-4o in two reference-free manners, which we refer to as “GRF” and “GEA”, respectively. The evaluation prompt of “GRF” borrows from Kocmi and Federmann (2023), and are illustrated in Appendix A. For “GEA”, the prompt mainly borrows from Wang et al. (2024), and we employ two variants of “GEA”, i.e., “GEA100” and “GEA5”. The evaluation prompt of “GEA100” is the same as the translation reward illustrated in § 2.1, while that of “GEA5” simply narrows the scoring scope of “GEA100” from a 100-point to a 5-point scale. Among them, “GRF” evaluates translations from a general perspective while “GEA5” and “GEA100” evaluate translation from a literary perspective. Since GRF, GEA5 and GEA100 need the costs of OpenAI’s API, we randomly select 400 samples from each test set (i.e., MetaphorTrans, O. Henry and Orbital), to conduct evaluation.
Backbones. We try to use LLMs with less than 10B parameters as the backbone of DeepTrans, given the high computation costs during RL training. Among all open-source LLMs, Qwen2.5-7B Yang et al. (2024) and Llama3-8B Grattafiori et al. (2024) are state-of-the-art choices. As pointed by Wang et al. (2024), Qwen2.5-7B-Instruct outperforms Llama3-8B-Instruct in literature translation. Therefore, we finally use Qwen2.5-7B-Instruct as the backbone of DeepTrans.
@@
Table 1: Comparisons between baselines and DeepTrans. Para.: Parameter; Reason.: Reasoning. Ckpt.: Checkpoint. “Training on MetaphorTrans” indicates whether the model is trained on the MetaphorTrans dataset.
@@
Table 2: Comparison results of DeepTrans and open-source baselines (< 30B parameters). The bold and the underline denote the best and second-best performances, respectively. $\dagger$” and“$\ddagger$” denote statistically significant better than the DRT-14B with t-test p < 0.01 and 0.05, respectively. $\diamondsuit$” denotes models are trained on MetaphorTrans.
@@
Table 3: Comparison results of DeepTrans and baselines (> 30B parameters or commercial LLMs). The bold and the underline denote the best and second-best performances, respectively.
@@ 3.2 Baselines
(1) General Non-reasoning LLMs. We use Llama-3.1-8B-Instruct Grattafiori et al. (2024), Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct Yang et al. (2024) and GPT-4o OpenAI (2024a) as baselines.
(2) General Reasoning LLMs. Marco-o1-7B Zhao et al. (2024), QwQ-32B-preview, QwQ-32B Qwen (2024), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1 Guo et al. (2025) and o1-preview OpenAI (2024b) are used as baselines.
@@
(4) MT Reasoning LLMs. Wang et al. (2024) introduce DRT-7B, DRT-8B and DRT-14B models, which are fine-tuned on the whole MetaphorTrans training data (with thought). Given sentences, these LLMs could first think about how to translate them and then provide the corresponding translations.
@@
3.3 Implementation Details.
Cold Start SFT. We randomly select 4K English sentences from WMT24555https://www2.statmt.org/wmt24/index.html in the general domain. Then, DeepSeek-R1 is employed to translate these sentences from English to Chinese in a deep reasoning manner. The synthesized 4K samples are further used to SFT DeepTrans (i.e., cold-start SFT). In detail, Llama-Factory framework Zheng et al. (2024) is used during the SFT stage. We conduct experiments on 8$\times$×NVIDIA H20 GPUs (96G) with 1e-5 learning rate and 8 (8$\times$1) batch size. DeepSpeed ZeRO-3 optimization Rasley et al. (2020) is also used during SFT. We set the number of SFT epochs to 2, and it costs about 1 GPU hour.
RL Training. We use GRPO RL algorithm implemented by verl [^6] https://github.com/volcengine/verl. During training, 2$\times$8 H20 GPUs are used, where 8 GPUs are used to deploy DeepSeek-v3 (awq quantization) as the reward model, and another 8 GPUs are used to optimize the policy model. We set the batch size to 64, the learning rate to 1e-6, the rollout number to 8 and the rollout temperature to 0.6, and the KL loss coefficient to 1e-3. The number of training epochs is set to 2. The RL training costs 2K GPU hours.
(a) Changes in overall rewards
(b) Changes in thought length
(c) Changes in CometKiwi
(d) Changes in GRF
(e) Changes in GEA5
(f) Changes in GEA100
Figure 2: Performance changes of DeepTrans-7B during RL training. Among them, (b), (d), (e) and (f) are conducted on the MetaphorTrans test set, while (c) is conducted on all three test sets. The training process totally involves 2 epochs and each epoch involves 300 training steps.
@@ 3.4 Main Results
As shown in Table 2, the results demonstrate the effectiveness of DeepTrans-7B. Specifically, it outperforms baselines with less than 7B parameters in terms of all metrics. Compared with all baselines (< 30B parameters), DeepTrans-7B achieves state-of-the-art performance in most metrics, especially GEA5 and GRF. Specifically, DeepTrans-7B outperforms its backbone (Qwen2.5-7B-Instruct) by 16.3%, 13.8% and 9.0% in terms of GEA5, GEA100 and GRF on the MetaphorTrans test set. As for strong baselines that are fine-tuned on MetaphorTrans (denoted with $\diamondsuit$”), they use both the source sentences and the synthesized translation targets. In contrast, DeepTrans-7B only uses the source sentences and achieves promising results. It is worth noting that training with RL does not need the synthesized translations (and thought process), and thus avoids the data quality issue as well as potential bias in synthesized data.
We also compare DeepTrans-7B with baselines (> 30B parameters or commercial LLMs), the results are shown in Table 3. DeepTrans-7B generally outperforms QwQ-32B-preview and DeepSeek-Qwen-32B, and achieves competitive results with GPT-4o, o1-preview, QwQ-32B and DeepSeek-R1.
@@
The man from the West unfolded the little piece of paper handed him. His hand was steady when he began to read, but it trembled a little by the time he had finished. The note was rather short. {CJK}UTF8gbsn 西方人展开递给他的一小张纸条。他开始阅读时手很稳,但读完后微微发抖。那张纸条很短。 [CometKiwi: 80.10] {CJK}UTF8gbsn 首先,西边来的男人展开递给他的一小张纸。他开始阅读时双手稳当,但读完后手指已微微颤抖。那张便条颇为简短。 [CometKiwi: 75.89] {CJK}UTF8gbsn 那位西部来的人接过那张折叠的小纸片,启口之际,双手仍显镇定,及至读罢,指尖却微微颤动。此笺虽属简短,却已道尽千般情愫[CometKiwi: 76.63] {CJK}UTF8gbsn 这位来自西方的男子,缓缓展开了递到他手中的小小纸片,当他开始阅读时,双手依然镇定,然而等到读罢之时,指尖已微微颤动。那张便条颇为简短。 [CometKiwi: 82.32] {CJK}UTF8gbsn 来自西方的男子,缓缓展开了递到他手中的小小纸片,起初,他的手显得十分稳当,可等到读罢,指尖竟微微颤抖起来。这封便条其实颇为简短。 [CometKiwi: 83.42]
When it was light enough Johnsy, the merciless, commanded that the shade be raised. {CJK}UTF8gbsn 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] {CJK}UTF8gbsn 当光线足够明亮时,那个无情的约翰西便下令拉起窗帘。 [CometKiwi: 81.56] {CJK}UTF8gbsn 待天色稍霁,约翰西这颗冷酷的心,便命将帘幕徐徐卷起[CometKiwi: 57.61] {CJK}UTF8gbsn 等到光线足够了,这个冷酷无情的约翰西,便下令将那阴影缓缓卷起。 [CometKiwi: 72.48] {CJK}UTF8gbsn 约翰西,这个冷酷无情的人,等到天色足够明亮,便命令将帘幕缓缓拉起。 [CometKiwi: 77.28]
When one loves one’s Art no service seems too hard. {CJK}UTF8gbsn 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] {CJK}UTF8gbsn 一旦钟爱艺术,便觉任何奉献皆非难事。 [CometKiwi: 81.86] {CJK}UTF8gbsn 一旦钟爱自家的艺术,任是何般劳顿艰辛,也觉微不足道,甘之如饴。 [CometKiwi: 65.53] {CJK}UTF8gbsn 只要钟爱自己的艺术,再大的付出也觉甘愿。 [CometKiwi: 67.99] {CJK}UTF8gbsn 只要钟爱自己的艺术,再大的付出也绝不嫌苦。 [CometKiwi: 76.16]
For two or three days she went out canvassing for pupils. One evening she came home elated. {CJK}UTF8gbsn 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] {CJK}UTF8gbsn 连续两三天,她外出物色学生。某晚归来时,她显得欣喜若狂。 [CometKiwi: 84.96] {CJK}UTF8gbsn 接连三四日,她外出寻觅门徒踪迹,直至某晚归来,已得诸多佳音满怀欣喜。 [CometKiwi: 64.59] {CJK}UTF8gbsn 连续一两天,她便出门去物色学生。一天傍晚,她兴冲冲地回到了家中。 [CometKiwi: 82.87] {CJK}UTF8gbsn 连续两三天,她外出寻觅学生上门求教。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 81.21]
@@ 3.5 Intermediate-Stage Analyses
To provide a deeper analysis of DeepTrans-7B, we further discuss the performance changes during RL training. Figure 2 shows the corresponding details, which we analyze from the following aspects:
@@
In terms of overall rewards, the training rewards and validation rewards generally increase along with the training process. The full score of the overall reward is 140 (c.f. Eq. 4), and DeepTrans-7B finally reaches about 120, which means $r_{\text{trans}}$ reaches at least 80 (c.f. Eq. 3). According to the definition of $r_{\text{trans}}$, 80 points indicate the model translations are between “very good translation” and “excellent translation”, showing the significant improvement brought by RL training. Similarly, as shown in Figure 2 (d), (e) and (f), model performance (in terms of GRF, GEA5 and GEA100) also generally increases along with the training process.
@@
In terms of thought length, the average length first significantly decreases from 800+ to 100 tokens in the first 100 steps, and then slowly increases to 200 tokens. This finding indicates that during RL training, the model will first focus on the translation results, and try to provide translations approved by the reward model. Meanwhile, the model starts to simplify the thought content, since no additional thought rewards are given. However, when the thought length is below a threshold, the model cannot provide a meaningful thought process with short words, and consequently, it will be punished by $r_{\text{thought}}$. Therefore, the model will stop simplifying the thought content, and learn to generate better translations and keep high-quality thought processes simultaneously. Later, during RL, the model will find a deeper thought that will lead to a better translation, and increase its thought length.
@@
In terms of CometKiwi, we find an interesting phenomenon: the performance drops a lot at first, and finally returns to the initial performance. To understand the reason behind this, we provide several cases generated by the intermediate-stage DeepTrans models in Table 4. We find that the model trained with 300 steps has a special characteristic: its translations are in strong Chinese classical literary style (e.g., ancient poetry). We mark several classical-style terms in purple, and these terms are also not commonly used in modern Chinese. The classical style might not be well captured by the CometKiwi calculating model [^7] https://huggingface.co/Unbabel/wmt22-cometkiwi-da, thus resulting in a low CometKiwi score. However, it still can be recognized by the GPT-4o evaluators. As shown in Figure 2, DeepTrans (300 steps) still outforms DeepTrans (cold start) in terms of GRF, GEA5 and GEA100. For this stylization phenomenon, we also want to discuss the following questions:
(1) Is stylization better or not? Though classical-style translations will provide some interesting results, we think this might lose the generalizability of the MT models. The classical-style translations are not suitable for all genres in literature translation. Our goal is to make the model provide translations that should be more easily absorbed by native people in target languages. The classical-style translations do not follow this goal, and it should be considered as a bias or preference. Therefore, we recommend that future work should also evaluate the stylization phenomenon during RL training.
(2) Why does DeepTrans first drop into the stylization phenomenon and finally get out of it? First, the model recognizes that the classical-style translations will receive a high translation reward (i.e., $r_{\text{trans}}$) during RL training, and thus, the model starts to provide classical-style translations. This also indicates the bias/preference of the reward model (i.e., DeepSeek-v3), which thinks classical-style translations are easy to understand for Chinese. However, such a stylization strategy will trap the model into a local optimal solution. Fortunately, the model discovers other better solutions during RL training, and finally discards the stylization strategy. Another important question is naturally raised: which factors lead the model to get out of the local optimal solution? We discuss it in § 4.3.
Figure 3: The performance comparisons between DeepTrans ($r_{\text{all}}$) and DeepTrans ($r_{\text{all (w/ lp)}}$) on MetaphorTrans test set.
@@
Figure 4: The performance comparisons among models trained with translation rewards at different scoring scales. The results are evaluated on the MetaphorTrans test set.

### v2

@@ 4.1 Experimental Setups
Data. We use MetaphorTrans Wang et al. (2025) in experiments, which is an English-Chinese literature MT data involving 19K training, 1K validation and 2K test samples. Each sample involves a source English sentence, the corresponding Chinese translation, and the thought process during translation. The source sentences are selected from English literature books, and generally contain metaphors or similes. The thought processes and translation results are synthesized via Qwen2.5-72B-Instruct Yang et al. (2024). We only use the source sentences during RL training.
In addition to the MetaphorTrans test set, we purchase electronic copies of two complete literature books, and use them to evaluate deep reasoning MT models: (1) The Essential O. Henry Collection (by O. Henry) and (2) Orbital (by Samantha Harvey). [^2] We will release the test data of The Essential O. Henry Collection, and release the data processing script of Orbital (the content of Orbital cannot be directly released due to copyright protection). Both books are rich in literary nuance, making it challenging for even human translators to achieve a free translation.
Metrics. Since (1) the references in MetaphorTrans are synthesized via LLMs, and are not verified by human translators; and (2) the golden references of The Essential O. Henry Collection and Orbital are missing, we adopt reference-free metrics in our experiments. Specifically, we use CometKiwi Rei et al. (2022) to evaluate the model translations, which judges whether a translation conveys the semantics of the source sentence. Moreover, following Wang et al. (2025), we use evaluators implemented using GPT-4o in two reference-free manners, which we refer to as GRF and GEA, respectively. The evaluation prompt of GRF borrows from Kocmi and Federmann (2023). [^3] Please refer to Figure 1 in Kocmi and Federmann (2023) For GEA, the prompt mainly borrows from Wang et al. (2025), and we employ two variants of GEA, i.e., GEA100 and GEA5. The evaluation prompt of GEA100 is the same as the translation reward illustrated in § 3.1, while that of GEA5 simply narrows the scoring scope of GEA100 from a 100-point to a 5-point scale. Among them, GRF evaluates translations from a general perspective while GEA5 and GEA100 evaluate translations from a literary perspective. The effectiveness of GRF in general translation and GEA in literary translation is demonstrated by Kocmi and Federmann (2023) and Wang et al. (2025), respectively. Since GRF, GEA5 and GEA100 need the costs of OpenAI’s API, we randomly select 400 samples from each test set to conduct evaluation.
Backbone. Given the high computation costs in RL, we try to use LLMs (< 10B parameters) as the backbone. Among all LLMs, Qwen2.5-7B Yang et al. (2024) and Llama3-8B Grattafiori et al. (2024) are state-of-the-art choices. Wang et al. (2025) show that Qwen2.5-7B outperforms Llama3-8B in literature translation. Thus, we use Qwen2.5-7B as the backbone of DeepTrans.
@@
Table 1: Comparisons between baselines. Para.: Parameter; Reason.: Reasoning. Ckpt.: Checkpoint.
@@
[…]
4.2 Implementation Details.
[…]
@@
Table 3: Comparison results of DeepTrans and baselines (> 30B parameters or commercial LLMs).
@@ 4.3 Baselines
(1) General Non-reasoning LLMs. We leverage Llama-3.1-8B-Instruct [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct Grattafiori et al. (2024), Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Yang et al. (2024) and GPT-4o OpenAI (2024a) as baselines.
(2) General Reasoning LLMs. QwQ-32B-preview [^9] https://huggingface.co/Qwen/QwQ-32B-Preview, QwQ-32B [^10] https://huggingface.co/Qwen/QwQ-32B Qwen (2024), Marco-o1-7B [^11] https://huggingface.co/AIDC-AI/Marco-o1 Zhao et al. (2024), DeepSeek-Qwen-7B [^12] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B, DeepSeek-Llama-8B [^13] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B, DeepSeek-Qwen-14B [^14] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-14B, DeepSeek-Qwen-32B [^15] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1161616https://huggingface.co/deepseek-ai/DeepSeek-R1 Guo et al. (2025) and o1-preview OpenAI (2024b) are used as baselines.
@@
(4) MT Reasoning LLMs. Wang et al. (2025) introduce DRT-7B [^17] https://huggingface.co/Krystalan/DRT-7B, DRT-8B [^18] https://huggingface.co/Krystalan/DRT-8B and DRT-14B [^19] https://huggingface.co/Krystalan/DRT-14B models, which are fine-tuned on the whole MetaphorTrans training data. Given sentences, these LLMs could first reason and then translate.
@@ 4.4 Main Results
As shown in Table 2, the experimental results verify the effectiveness of DeepTrans-7B. Specifically, compared with all baselines (< 30B parameters), DeepTrans-7B achieves state-of-the-art performance in most metrics, especially GEA5 and GRF. DeepTrans-7B outperforms its backbone (Qwen2.5-7B-Instruct) by 16.3%, 13.8% and 9.0% in terms of GEA5, GEA100 and GRF on MetaphorTrans. As for strong baselines that are fine-tuned on MetaphorTrans (denoted with $\diamondsuit$”), they use the source sentences and the synthesized literature translations. In contrast, DeepTrans-7B only uses the source sentences and achieves promising results, and thus avoids the data quality issue and potential bias in synthesized data.
To figure out the effects of the cold start SFT and the RL training, we also compare DeepTrans-7B with its two variants: (1) DeepTrans-7B (Cold Start) is only trained via the cold start SFT and does not incorporate RL training. In contrast, (2) DeepTrans-7B (Direct RL) is trained solely using RL training without the cold start SFT. The experimental results in Table 2 show that both variants underperform the original DeepTrans-7B across all metrics, demonstrating the rationality and the effectiveness of the two-stage training process. The cold start SFT lets the backbone LLM (i.e., Qwen2.5-7B-Instruct) quickly learn the deep reasoning translation task, and brings MT performance improvements to the model. Without the cold start SFT, DeepTrans-7B (Direct RL) only achieves sub-optimal performance. After the cold start SFT, the RL training stage further guides the model to acquire effective translation strategies. The substantial improvements of DeepTrans-7B compared to DeepTrans-7B (Cold Start) verify the superiority of the RL training.
Model Flu. Sem. Lit.
MetaphorTrans
Qwen2.5-14B-SFT 0.044 -0.009 -0.105
DRT-14B 0.079 0.079 0.109
DeepTrans-7B (Cold Start) -0.258 -0.291 -0.269
DeepTrans-7B 0.135 0.221 0.265
O. Henry
Qwen2.5-14B-SFT -0.020 -0.019 -0.170
DRT-14B 0.088 0.072 0.129
DeepTrans-7B (Cold Start) -0.245 -0.294 -0.241
DeepTrans-7B 0.177 0.241 0.282
Orbital
Qwen2.5-14B-SFT -0.010 0.001 -0.120
DRT-14B 0.091 0.082 0.109
DeepTrans-7B (Cold Start) -0.238 -0.274 -0.241
DeepTrans-7B 0.157 0.191 0.252
Table 4: Human evaluation results in terms of fluency, semantic accuracy and literary quality.
We also compare DeepTrans-7B with baselines (> 30B or commercial LLMs). As shown in Table 3, DeepTrans-7B generally outperforms QwQ-32B-preview and DeepSeek-Qwen-32B, and achieves competitive results with QwQ-32B, DeepSeek-R1, GPT-4o and o1-preview. This result demonstrates the superiority of DeepTrans-7B, with only 7B parameters, which shows strong performance in deep reasoning MT.
(a) Overall rewards ($r_{\text{all}}$)
(b) Thought length (token-level)
(c) CometKiwi
(d) GRF
(e) GEA5
(f) GEA100
Figure 2: Performance changes of DeepTrans-7B during RL training. Among them, (c) is conducted on all three test sets, while others are conducted on MetaphorTrans. The horizontal axis denotes the number of training steps, and there are 600 steps (2 epochs) in total. The vertical axis denotes the value of the corresponding metrics.
@@
When it was light enough Johnsy, the merciless, commanded that the shade be raised. 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] 当光线足够明亮时,那个无情的约翰西便下令拉起窗帘。 [CometKiwi: 81.56] 待天色稍霁,约翰西这颗冷酷的心,便命将帘幕徐徐卷起[CometKiwi: 57.61] 等到光线足够了,这个冷酷无情的约翰西,便下令将那阴影缓缓卷起。 [CometKiwi: 72.48] 约翰西,这个冷酷无情的人,等到天色足够明亮,便命令将帘幕缓缓拉起。 [CometKiwi: 77.28]
When one loves one’s Art no service seems too hard. 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] 一旦钟爱艺术,便觉任何奉献皆非难事。 [CometKiwi: 81.86] 一旦钟爱自家的艺术,任是何般劳顿艰辛,也觉微不足道,甘之如饴。 [CometKiwi: 65.53] 只要钟爱自己的艺术,再大的付出也觉甘愿。 [CometKiwi: 67.99] 只要钟爱自己的艺术,再大的付出也绝不嫌苦。 [CometKiwi: 76.16]
For two or three days she went out canvassing for pupils. One evening she came home elated. 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] 连续两三天,她外出物色学生。某晚归来时,她显得欣喜若狂。 [CometKiwi: 84.96] 接连三四日,她外出寻觅门徒踪迹,直至某晚归来,已得诸多佳音满怀欣喜。 [CometKiwi: 64.59] 连续一两天,她便出门去物色学生。一天傍晚,她兴冲冲地回到了家中。 [CometKiwi: 82.87] 连续两三天,她外出寻觅学生上门求教。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 81.21]
@@
4.5 Human Evaluation
We conduct human evaluation to further evaluate the performance of DeepTrans-7B, DeepTrans-7B (Cold Start), DRT-14B and Qwen2.5-14B-SFT. We randomly select 200 samples from each test set, and employ five human evaluators with high levels of fluency in English and Chinese to assess the generated translations. The human evaluation focuses on three key aspects: fluency (Flu.), semantic accuracy (Sem.) and literary quality (Lit.). Following Kiritchenko and Mohammad (2017) and Wang et al. (2025), evaluators are tasked with identifying the best and worst translations for each aspect. The result scores are calculated based on the percentage of times each model is selected as best minus the times it is selected as worst. Consequently, the final scores range from -1 (indicating the worst performance) to 1 (indicating the best performance). As shown in Table 4, DeepTrans-7B significantly outperforms the others, especially in literary quality. This result demonstrates the superiority and effectiveness of DeepTrans-7B. The Fleiss’ Kappa scores Fleiss (1971) of Flu., Sem. and Lit. are 0.68, 0.70 and 0.74, respectively, indicating a good inter-agreement among evaluators.
@@ 4.6 Intermediate-Stage Analyses
To provide a deeper analysis of DeepTrans, we discuss the performance changes during RL training. Figure 2 shows the corresponding details, which we analyze from the following aspects:
@@
In terms of overall rewards (c.f. Figure 2 (a)), the training rewards and validation rewards generally increase along with the training process. The full score of the overall reward is 140 (according to Eq. 4), and DeepTrans-7B finally reaches about 120, which means $r_{\text{trans}}$ reaches at least 80. According to the definition of $r_{\text{trans}}$, 80 points indicate the translations are between “very good translation” and “excellent translation”, showing the superiority of DeepTrans-7B. Similarly, as shown in Figure 2 (d), (e) and (f), model performance in GRF, GEA5 and GEA100 also generally increases along with the training process. After RL training, DeepTrans-7B demonstrates a substantial performance enhancement compared to the initial model, which is only trained via the cold start SFT. This also verifies the effectiveness of RL training.
@@
In terms of thought length (c.f. Figure 2 (b)), the average length first significantly decreases from 800+ to 100 tokens, and then slowly increases to 200 tokens. This finding indicates that DeepTrans-7B first generates plenty of thought content owing to the cold start SFT, and thus can obtain high thought rewards. During RL training, the model first preliminarily focuses on the translation results instead of the thought content, and it starts to simplify the thought content, since no additional thought rewards could be given. However, when the thought length is below a threshold, the model cannot provide a meaningful thought content with short words, and it will be punished by $r_{\text{thought}}$. Then, the model will stop simplifying the thought content, and learn to generate better translations and keep high-quality thought processes simultaneously. Later, along with RL training, the model will find deeper thoughts that will lead to better translations, and learn to improve its thinking.
@@
In terms of CometKiwi (c.f. Figure 2 (c)), we find an interesting phenomenon: the performance drops a lot at first, and finally returns to the initial performance. To understand the reason behind this, we provide several cases generated by the intermediate-stage DeepTrans-series models. As shown in Table 5, we find that the model trained with 300 steps (i.e., 1 epoch) has a special characteristic: its translations exhibit a distinctly classical Chinese literary style (e.g., ancient poetry). We mark several classical-style terms in purple, and these terms are also not commonly used in modern Chinese. The classical style might not be well captured by the CometKiwi. However, it still can be recognized by the GPT-4o evaluators. As shown in Figure 2, DeepTrans (300 steps) still outperforms DeepTrans (cold start) in terms of GRF, GEA5 and GEA100. For this stylization phenomenon, we also want to discuss the following questions: (1) Is stylization better or not? Though classical-style translations will provide some interesting results, we think this might lose the generalizability of the MT models. The classical-style translations are not suitable for all genres in literature. Our goal is to make the model provide translations that should be more easily absorbed by native people in target languages. The classical-style translations do not follow this goal, and it should be considered a bias or preference. Thus, we recommend that future work should also evaluate the stylization phenomenon during RL training. (2) Why does DeepTrans first drop into the stylization phenomenon and finally get out of it? First, the model recognizes that the classical-style translations will receive a high translation reward, and thus, the model starts to provide classical-style translations. This also indicates a bias/preference of the reward model. However, such a stylization strategy will trap the model into a local optimal solution. Fortunately, the model discovers other better solutions during exploration in RL, and finally discards the stylization strategy. Another important question is naturally raised: which factors lead the model to get out of the local optimal solution? We will discuss it in § 5.3.
Figure 3: The comparisons between DeepTrans and DeepTrans (w/ LP) on MetaphorTrans.
@@
Figure 4: Comparisons of models trained with translation rewards across different scoring scales. The results are evaluated on the MetaphorTrans test set.